A battery powered torch uses a 9 V battery and dissipates 0.25 W of electrical power in the bulb.

If the torch is used for 1.0 hours then how much charge has passed through the bulb? (to 2 s.f and in C)


1
Expert's answer
2021-05-25T10:10:31-0400

"P = 0.25 \;W \\

V = 9 \;V \\

t = 1.0 \;hour = 1 \times 60 \times 60 = 3600 \;sec \\

I = \frac{P}{V} \\

= \frac{0.25}{9}=0.02777 \;A \\

Q = It \\

= 0.02777 \times 3600 = 100 \;C"

Answer: 100 C